Rhodes Corporation: Income Statements for Year Ending December 31 (Millions of Dollars)

2016 2015
Sales 2875.0 2500.0
Operating costs excluding 2372.0 2125.0
depreciation
Depreciation and amortization 69.0 55.0
Earnings before interest and taxes 434.0 320.0
Less: Interest 62.0. 54.0
Pre-tax income 372.0 266.0
Taxes (40%). 148.8 106.4
Net Income available to common 223.2 159.6
stockholders
Common dividends 201.0. 128.0
Rhodes Corporation Balance Sheet as of December 31 (Millions of Dollars)

Assets
Cash 40.0 35.0
Short-Term investments 15.0. 13.0
Accounts Receivable 303.0. 275.0
Inventories 520.0. 400.0
Total current assets 878.0. 723.0
Net Plant and equipment 688.0 550.0
Total Assets 1566.0. 1273.0

Liabilities and Equity
Accounts Payable 288.0 250.0
Accruals 180.0 150.0
Notes Payable 58.0 50.0
Total current liabilities 526.0 450.0
Long Term debt 575.0 500.0
Total liabilities. 1101.0 950.0
Common Stock 395.8 276.0
Retained Earnings 69.2 47.0
Total common equity 465.0 323.0
Total liabilities and equity 1566.0 1273.0


a.What is the net operating profit after taxes (NOPAT) for 2016?
b. What are the amounts of net operating working capital for both years?
2016
2015
c. What are the amounts of total net operating capital for both years?
2016
2015
d. What is the free cash flow for 2016?
e. What is the ROIC for 2016?
f. How much of the FCF did Rhodes use for each of the following purposes: after tax interest: after tax interest, net debt repayment, dividends, net stock repurchases, and net purchases of short term investments?
After tax interest payment
Reduction in debt
Payments of dividends
Repurchase stock
Purchase of short term investment
Answer #1

IT is just missing part f. Hope this helps you

a. The NOPAT for 2016 is 148.8 millions

b. The formula to get the Net Operating Working Capital is
= Operating Current Assets − Operating Current Liabilities.

(Cash+ Accounts Receivable + Inventories)

For 2016 For 2015

  • Cash 40.0. 35.0
  • Accounts receivable 303.0. 275.0  
  • Inventories 520.0. 400.0
  • Operating current assets=863.0. 710

Operating current Liabilities

(Accounts Payable + Accrued Expenses).

For 2016. For 2015

Accounts payable 288.0. 250

Accrued expenses180. 150

Notes payable 58. 50

Op. current liabilities 526. 450

Net operating working capital

For 2016. For 2015   

863.0- 526=337   710- 450= 260

.c. Total net operating capital.

Total Net Operating Capital = Net Operating Working Capital + Non-current Operating Assets

For 2016 For 2015

337 + 688=1,025. 260 +550= 810

d. Free cash flow

Net income+ depreciation and amortization (+or-) the change is working capital

223.2+69+142= 434.2

e. ROIC.

(net income - dividend) / (debt + equity)=

(223.2- 201)/1566=0.0141762452=1.42